pre

Hoose af constan value is a state Rintion Heense y has a uiqne value even though ntate Heeause at constant voluime AUW0, le a ate fton Casulaie the diffrenee between the molar internal eneray and the mo s peon h 24700 kJ/mol 35.859 JAnol d I8ON Amol 8.3145 Jnol
Answer #1

10) Calculate the difference between the molar internal energy and the molar enthalpy for a perfect gas at 298.15K

For a perfect gas, we can use the formula:

H - U = RT

Where H is the molar enthalpy and U is the molar internal energy. R = 8.314 J/molK and T = 298.15K

We are looking for H - U, so just replace the values given above in the formula:

H - U = (8.314 J/molK)(298.15K)

H - U = 2478.8 J/mol

H - U = 2.479 kJ/mol

So, the correct answer is b) 2.4790 kJ/mol